# 3D-GP-LMVIC: Learning-based Multi-View Image Coding with 3D Gaussian Geometric Priors

### v1

@@ 5 Conclusion
In this paper, we present 3D-GP-LMVIC, a novel learning-based multi-view image coding framework incorporating 3D Gaussian geometric priors. This framework exploits these geometric priors to estimate complex disparities and masks between views for effectively utilizing reference view information in the compression process. Additionally, we propose a depth map compression model designed to compactly and accurately represent the geometry of each view, incorporating a cross-view depth prediction module to capture inter-view geometric correlations. Moreover, we introduce a multi-view sequence ordering method for unordered sequences, enhancing the overlap between adjacent views by defining an inter-view distance measure to guide the sequence ordering. Experimental results confirm that 3D-GP-LMVIC surpasses existing learning-based coding schemes in compression efficiency while maintaining competitive coding speed.
Appendix
Figure 6: The proposed depth map compression model. Convolution and deconvolution layers are specified by the tuple (number of output channels, kernel size, stride). ’Q’ denotes the quantization step. ’AE’ and ’AD’ stand for arithmetic encoder and arithmetic decoder, respectively.
Appendix A Supplementary Information for the Depth Map Compression Model
@@
As illustrated in Figure 6, during the compression and decompression of $\bm{d}_{n}$, $\bm{\hat{d}}_{n-1}$ is initially processed by the depth prediction extractor, which extracts multi-scale depth prediction features and associated feature masks. These extracted features and masks are then integrated into the depth backbone encoder and decoder via the depth context integration (DCI) module. Detailed explanations of the depth prediction extractor and the DCI module are provided in the subsequent content.
@@
Depth prediction extractor. As illustrated by the brown dotted-dashed line box in Figure 6, we first utilize the proposed cross-view depth prediction (CVDP) module to predict the depth map $\bm{d}_{n,\text{p}}\in\mathbb{R}^{W\times H}$ and the associated mask $\bm{d}_{n,\text{m}}\in\mathbb{R}^{W\times H}$ for the $n$-th view, based on $\bm{\hat{d}}_{n-1}$. Specifically, for each pixel $(x_{n-1},y_{n-1})$ in the $(n-1)$-th view and its corresponding reconstructed depth $\hat{d}_{n-1}$, the CVDP module determines the corresponding pixel coordinates $(x_{n},y_{n})$ and the depth prediction $d^{\prime}_{n}$ in the $n$-th view using the method described in Eq. 3. The depth at the nearest grid point $(\lfloor x_{n}-0.5\rceil,\lfloor y_{n}-0.5\rceil)$ is then set to $d^{\prime}_{n}$:
$\bm{d}_{n,\text{p}}[\lfloor x_{n}-0.5\rceil,\lfloor y_{n}-0.5\rceil]=d^{\prime}_{n}.$ (9)
@@
This cross-view depth prediction is applied to each pixel in the $(n-1)$-th view to construct $\bm{d}_{n,\text{p}}$. If multiple pixel coordinates map to the same grid point, the depth prediction for that point is set to the minimum of these predicted depths. Additionally, if a grid point has no corresponding pixel coordinates, its depth prediction value is set to 0. The mask $\bm{d}_{n,\text{m}}$ indicates whether each grid point has at least one corresponding pixel coordinate, with values set to 1 where a correspondence exists and 0 otherwise.
@@
Subsequently, $\bm{d}_{n,\text{p}}$ is input into the depth prediction feature extractor to obtain multi-scale depth prediction features $\{\bm{g}^{i}_{n,\text{p}}\mid i=1,2,3,4\}$. Concurrently, the mask $\bm{d}_{n,\text{m}}$ is processed by the depth mask extractor to derive the associated multi-scale feature masks $\{\bm{g}^{i}_{n,\text{m}}\mid i=1,2,3,4\}$.
@@
Depth Context Integration Module. Each DCI module integrates the input features $\bm{g}^{i^{*}}_{n}$ from the backbone network with $\bm{g}^{i}_{n,\text{p}}$ through channel-wise concatenation, followed by element-wise multiplication with $\bm{g}^{i}_{n,\text{m}}$ to produce the output feature $\bm{g}^{i}_{n}$:
$\bm{g}^{i}_{n}=(\bm{g}^{i^{*}}_{n}\oplus\bm{g}^{i}_{n,\text{p}})\odot\bm{g}^{i}_{n,\text{m}},$ (10)
@@
Appendix B Proof of $\mathcal{D}_{\mathcal{V}}(i,j)$ as a Distance Measure for 2-Norm and Frobenius Norm
B.1 Proof for 2-Norm
B.1.1 Definition
Definition 1.
@@ B.1.1 Definition
B.1.2 Lemma
Lemma 1.
For any $u=(A,B)$ and $v=(C,D)$ as defined in Definition 1, the following inequality holds:
@@ B.1.3 Theorem
The second inequality follows from Lemma 1. ∎
B.2 Proof for Frobenius Norm
B.2.1 Definition
Definition 2.
For $u=(A,B)$ and $v=(C,D)$ as defined in Definition 1, we define $(u,v)_{F}=\mathrm{tr}\left(AC^{T}+BD^{T}\right)$.
B.2.2 Lemma
Lemma 2.
For $u$ and $v$ as defined in Definition 2, the following inequality holds:
@@ B.2.2 Lemma
The method of proof is analogous to that used in Lemma 1. By leveraging the properties of the trace and following a similar reasoning process, the result is derived. ∎
B.2.3 Theorem
Theorem 2.
@@ B.2.3 Theorem
Non-negativity: The proof follows a similar approach to that of Theorem 1, so we omit the details here.
@@
Triangle Inequality: For views $i$, $j$, and $k$, we follow the same definitions of $A_{i,j}$, $B_{i,j}$, $A_{j,k}$, $B_{j,k}$, $A_{k,i}$, $B_{k,i}$, $u_{j,k}$, and $u_{k,i}$ as in the proof of the triangle inequality in Theorem 1. Starting from the fourth equality in the proof of symmetry, we have:
@@
Appendix C Experimental Details
@@
Table 3: Average alignment quality (PSNR, MS-SSIM) and runtime of different alignment methods on the Train scene of the Tanks&Temples dataset.
Metrics Methods
HT PM SPyNet PWC-Net FlowFormer++ 3D-GP-A
PSNR 15.16 17.94 16.12 17.59 18.08 18.14
MS-SSIM 0.5435 0.7633 0.6289 0.7707 0.7863 0.8053
Runtime 43ms 207ms 7ms 11ms 836ms 14ms
Reference View
Target View
HT
12.82/0.6065
PWC-Net
9.65/0.2840
PM
10.05/0.4078
FlowFormer++
13.94/0.7216
SPyNet
10.25/0.1990
3D-GP-A
13.94/0.7217
@@
Figure 8: Visual examples of 3D-GP-A and the mask from Eq. 4.
Appendix D Alignment Experiments
To evaluate the effectiveness of the proposed 3D Gaussian geometric priors-based alignment method (3D-GP-A), we conducted alignment experiments on the Train scene from the Tanks&Temples dataset. The baselines for comparison include alignment methods used in learning-based multi-view image codecs: homography transformation (HT) (Deng et al., 2021), patch matching (PM) (Huang et al., 2023), and optical flow estimation methods: SPyNet (Ranjan & Black, 2017), PWC-Net (Sun et al., 2018), and FlowFormer++ (Shi et al., 2023). Alignment quality was assessed by computing PSNR and MS-SSIM between the aligned reference view images and the target view images. Table 3 summarizes the average alignment quality and runtime for each method. The proposed 3D-GP-A method outperformed the baselines in both PSNR and MS-SSIM, indicating its effectiveness in capturing complex disparities between views while maintaining competitive runtime performance. Figure 7 provides visual comparisons, demonstrating that 3D-GP-A achieves closer alignment with the target view images. However, in the magnified red box, two iron bars appear, while only the left bar exists in the target view. This ghosting effect occurs because the train’s outer shell behind the right bar in the target view is occluded in the reference view, causing misalignment. The mask $\bm{x}_{n,\text{m}}$ in Eq. 4 can indicate occluded regions. Figure 8 shows visual examples of 3D-GP-A along with the corresponding masks. Notably, ghosting artifacts due to occlusion, such as those involving the iron bars and the edge of the train shell, are effectively identified by the mask, aiding the codec in filtering out irrelevant information when merging features from the reference view.

### v2

@@ 5 Conclusion
In this paper, we present 3D-LMVIC, a novel learning-based multi-view image coding framework incorporating 3D Gaussian geometric priors. This framework exploits these geometric priors to estimate complex disparities and masks between views for effectively utilizing reference view information in the compression process. Additionally, we propose a depth map compression model designed to compactly and accurately represent the geometry of each view, incorporating a cross-view depth prediction module to capture inter-view geometric correlations. Moreover, we introduce a multi-view sequence ordering method for unordered sequences, enhancing the overlap between adjacent views by defining an inter-view distance measure to guide the sequence ordering. Experimental results confirm that 3D-LMVIC surpasses existing learning-based coding schemes in compression efficiency while achieving accurate disparity estimation.
Appendix A Disparity and Mask Estimation Algorithm
Algorithm 1 Disparity and Mask Estimation
  Input: Depth estimation function $GSDE$, intrinsic matrix $K$, extrinsic matrices $V_{n}$ and $V_{n-1}$
  Output: Disparity map $\boldsymbol{\Delta}_{n}$, mask $\boldsymbol{x}_{n,\text{m}}$
  $\boldsymbol{d}_{n}\leftarrow GSDE(K,V_{n})$
  $\boldsymbol{d}_{n-1}\leftarrow GSDE(K,V_{n-1})$
  $\boldsymbol{\Delta}_{n},\boldsymbol{d}^{\prime}_{n-1}\leftarrow\text{DisparityEstimation}(\boldsymbol{d}_{n},K,V_{n},V_{n-1})$
  $\boldsymbol{x}_{n,\text{m}}\leftarrow\text{MaskEstimation}(\boldsymbol{\Delta}_{n},\boldsymbol{d}^{\prime}_{n-1},\boldsymbol{d}_{n-1})$
Figure 8: The architecture of the proposed depth map compression model. ’LR’ represents the Leaky ReLU activation function, ’Q’ denotes the quantization operation, and ’AE’/’AD’ refer to the arithmetic encoder/decoder, respectively.
Appendix B Supplementary Information for the Depth Map Compression Model
@@
As illustrated in Figure 8, during the compression and decompression of $\boldsymbol{d}_{n}$, $\boldsymbol{\hat{d}}_{n-1}$ is initially processed by the depth prediction extractor, which extracts multi-scale depth prediction features and associated feature masks. These extracted features and masks are then integrated into the depth backbone encoder and decoder via the depth context integration (DCI) module. Detailed explanations of the depth prediction extractor and the DCI module are provided in the subsequent content.
@@
Depth prediction extractor. As illustrated in Figure 8, we first utilize the proposed cross-view depth prediction (CVDP) module to predict the depth map $\boldsymbol{d}_{n,\text{p}}\in\mathbb{R}^{W\times H}$ and the associated mask $\boldsymbol{d}_{n,\text{m}}\in\mathbb{R}^{W\times H}$ for the $n$-th view, based on $\boldsymbol{\hat{d}}_{n-1}$. Specifically, for each pixel $(x_{n-1},y_{n-1})$ in the $(n-1)$-th view and its corresponding reconstructed depth $\hat{d}_{n-1}$, the CVDP module determines the corresponding pixel coordinates $(x_{n},y_{n})$ and the depth prediction $d^{\prime}_{n}$ in the $n$-th view using the method described in (3). The depth at the nearest grid point $(\lfloor x_{n}-0.5\rceil,\lfloor y_{n}-0.5\rceil)$ is then set to $d^{\prime}_{n}$:
$\boldsymbol{d}_{n,\text{p}}[\lfloor x_{n}-0.5\rceil,\lfloor y_{n}-0.5\rceil]=d^{\prime}_{n}.$ (10)
@@
This cross-view depth prediction is applied to each pixel in the $(n-1)$-th view to construct $\boldsymbol{d}_{n,\text{p}}$. If multiple pixel coordinates map to the same grid point, the depth prediction for that point is set to the minimum of these predicted depths. Additionally, if a grid point has no corresponding pixel coordinates, its depth prediction value is set to 0. The mask $\boldsymbol{d}_{n,\text{m}}$ indicates whether each grid point has at least one corresponding pixel coordinate, with values set to 1 where a correspondence exists and 0 otherwise.
@@
Subsequently, $\boldsymbol{d}_{n,\text{p}}$ is fed into the depth prediction feature extractor to produce multi-scale depth prediction features, denoted as $\{\boldsymbol{g}^{i}_{n,\text{p}}\mid i=1,2,3,4\}$. Concurrently, the mask $\boldsymbol{d}_{n,\text{m}}$ is processed by the depth mask extractor to derive the associated multi-scale feature masks $\{\boldsymbol{g}^{i}_{n,\text{m}}\mid i=1,2,3,4\}$.
@@
Depth Context Integration Module. Each DCI module integrates the input features $\boldsymbol{g}^{i^{*}}_{n}$ from the backbone network with $\boldsymbol{g}^{i}_{n,\text{p}}$ through channel-wise concatenation, followed by element-wise multiplication with $\boldsymbol{g}^{i}_{n,\text{m}}$ to produce the output feature $\boldsymbol{g}^{i}_{n}$:
$\boldsymbol{g}^{i}_{n}=(\boldsymbol{g}^{i^{*}}_{n}\oplus\boldsymbol{g}^{i}_{n,\text{p}})\odot\boldsymbol{g}^{i}_{n,\text{m}},$ (11)
@@
Appendix C Proof of $\mathcal{D}_{\mathcal{V}}(i,j)$ as a Distance Measure for 2-Norm and Frobenius Norm
C.1 Proof for 2-Norm
C.1.1 Definition
Definition C.1.
@@ C.1.1 Definition
C.1.2 Lemma
Lemma C.2.
For any $u=(A,B)$ and $v=(C,D)$ as defined in Definition C.1, the following inequality holds:
@@ C.1.3 Theorem
The second inequality follows from Lemma C.2. ∎
C.2 Proof for Frobenius Norm
C.2.1 Definition
Definition C.4.
For $u=(A,B)$ and $v=(C,D)$ as defined in Definition C.1, we define $(u,v)_{F}=\mathrm{tr}\left(AC^{T}+BD^{T}\right)$.
C.2.2 Lemma
Lemma C.5.
For $u$ and $v$ as defined in Definition C.4, the following inequality holds:
@@ C.2.2 Lemma
The method of proof is analogous to that used in Lemma C.2. By leveraging the properties of the trace and following a similar reasoning process, the result is derived. ∎
C.2.3 Theorem
Theorem C.6.
@@ C.2.3 Theorem
Non-negativity: The proof follows a similar approach to that of Theorem C.3, so we omit the details here.
@@
Triangle Inequality: For views $i$, $j$, and $k$, we follow the same definitions of $A_{i,j}$, $B_{i,j}$, $A_{j,k}$, $B_{j,k}$, $A_{k,i}$, $B_{k,i}$, $u_{j,k}$, and $u_{k,i}$ as in the proof of the triangle inequality in Theorem C.3. Starting from the fourth equality in the proof of symmetry, we have:
@@
Appendix D Experimental Details
@@
Figure 9: Visual examples of proposed alignment method and the mask from (4).
Appendix E Supplementary Alignment Experiments
Figure 9 shows visual examples of proposed alignment method along with the corresponding masks. Notably, ghosting artifacts due to occlusion, such as those involving the iron bars and the edge of the train shell, are effectively identified by the mask, aiding the codec in filtering out irrelevant information when merging features from the reference view.
Ground truth
0.9482/36.88/0.9900
1.1425/36.06/0.9917
1.1475/34.95/0.9938
0.8110/37.85/0.9952
LDMIC
0.7685/36.18/0.9887
0.8934/35.34/0.9913
0.9431/33.52/0.9927
0.6574/36.75/0.9950
BiSIC
0.6050/37.73/0.9908
0.6545/37.14/0.9924
0.7596/36.72/0.9946
0.6063/38.89/0.9950
3D-LMVIC
Figure 10: Visual Comparison of LDMIC, BiSIC, and 3D-LMVIC on the Tanks&Temples Dataset. Compression performance is reported as bpp/PSNR/MS-SSIM.
Appendix F Visualization
In Figure 10, we present examples from the Tanks&Temples dataset to visually compare the performance of LDMIC, BiSIC, and 3D-LMVIC. The results demonstrate that 3D-LMVIC preserves more texture details and achieves higher reconstruction quality for elements like branches, humans, and text, while consuming fewer bits.
Table 3: Complexity of learning-based image codecs evaluated on images with the resolution as 978×546 in the Tanks&Temples dataset.
Codecs MACs Enc. MACs Dec. Params Enc. Params Dec. Time Enc. Time Dec. Memory
HESIC+ 48.16G 134.31G 17.18M 15.1M 4.35s 10.73s 2248M
MASIC 65.62G 511.34G 32.03M 30.73M 4.38s 10.78s 5202M
SASIC 91.80G 438.09G 3.57M 4.44M 0.06s 0.09s 4498M
LDMIC-Fast 37.49G 94.43G 7.73M 11.15M 0.11s 0.09s 1168M
LDMIC 30.91G 87.84G 7.73M 11.15M 4.24s 10.63s 1096M
BiSIC-Fast 1880G (Enc.+Dec.) 85.9M (Enc.+Dec.) - - 3552M
BiSIC 1770G (Enc.+Dec.) 78.21M (Enc.+Dec.) - - 3006M
3D-LMVIC 479.43G 436.16G 41.92M 36.87M 0.19s 0.18s 3164M
Appendix G Complexity Analysis
Table 3 summarizes the Multiply-Accumulate Operations (MACs), model parameters, coding speed, and memory usage of eight learning-based image codecs. These evaluations were conducted on a platform equipped with an Intel(R) Xeon(R) Gold 6330 CPU @ 2.00GHz and an NVIDIA RTX A6000 GPU. The neural network components were executed on the GPU, while entropy coding was performed on the CPU.
Due to the absence of separate encoder and decoder implementations in the open-source code of BiSIC, we measured only its overall computational complexity. The proposed 3D-LMVIC demonstrates computational complexity within an acceptable range, comparable to the SOTA BiSIC and slightly better than BiSIC-Fast. Specifically, 3D-LMVIC achieved encoding and decoding times of 0.19s and 0.18s, respectively, ranking it among the faster methods.
While the MACs of the 3D-LMVIC encoder are relatively high, they remain lower than those of BiSIC, which employs a symmetric encoder-decoder structure. For BiSIC, we estimate that the MACs for its encoder and decoder each account for approximately half of the total MACs. Additionally, the inclusion of a depth map codec in 3D-LMVIC contributes to the higher MACs and model parameter count.
Table 4: BDBR of 3D-LMVIC relative to HEVC.
Methods Tanks&Temples Mip-NeRF 360 Deep Blending
PSNR MS-SSIM PSNR MS-SSIM PSNR MS-SSIM
3D-LMVIC -20.69% -40.75% -14.48% -22.06% -17.29% -43.06%
Appendix H Supplementary Coding Performance
We present a supplementary comparison of the coding performance between the proposed 3D-LMVIC and the HEVC video coding standard. The multi-view sequences are treated as a single video and compressed using HEVC with the lowdelay_P configuration and YUV444 input format. HEVC’s coding efficiency is evaluated using the HM-18.0 software [^3] https://vcgit.hhi.fraunhofer.de/jvet/HM/-/tags. Table 4 reports the BDBR of 3D-LMVIC relative to HEVC. On the three datasets, 3D-LMVIC consistently surpasses HEVC in both PSNR and MS-SSIM, demonstrating its effectiveness in reducing inter-view redundancy in multi-view sequences.